# Internal symmetry to the rescue: well-posed 1+1 evolution of self-interacting vector fields

### v1

@@ I Introduction
These issues were also identified in the generalized Proca theory Tasinato (2014); Heisenberg (2014); Allys et al. (2016); Beltran Jimenez and Heisenberg (2016); Gallego Cadavid and Rodriguez (2019), a vector-tensor formulation inspired by Horndeski theory Horndeski (1974); Deffayet et al. (2011). Thus, it seems that this issue is not exclusive to Proca fields alone, it also arises in generalizations of the Proca theory Garcia-Saenz et al. (2021); Ünlütürk et al. (2023). These findings pose a challenge to the effective field theory description of massive vector fields, particularly those based on effective description of interactions, including dark matter scenarios Holdom (1986); Graham et al. (2016); Agrawal et al. (2020); Goodsell et al. (2009); Bastero-Gil et al. (2019); Ema et al. (2019); Kolb and Long (2021); Ahmed et al. (2020); Caputo et al. (2021); Antypas et al. (2022). A potential resolution to this issue is that the effective field theory description breaks down before the ghost instability develops Aoki and Minamitsuji (2022). Another approach consists in fixing the equations Barausse et al. (2022); Rubio et al. (2024), a procedure that “correct” the hyperbolicity character of the system Cayuso et al. (2017). See also Ref. Coates and Ramazanoğlu (2023c) for a discussion on the shortcomings of this approach.
@@
At this point, the reader may wonder why we focus specifically on non-Abelian SU(2) fields. The motivation stems from their potential to drive various phenomena in both cosmology Maleknejad and Sheikh-Jabbari (2013, 2011); Adshead and Wyman (2012); Garnica et al. (2022); Guarnizo et al. (2020); Garcia-Serna et al. (2025) and astrophysics Bartnik and Mckinnon (1988); Bizon (1990); Greene et al. (1993); Dzhunushaliev and Folomeev (2020); da Rocha (2020); Jain (2022); Martinez et al. (2023); Gómez and Rodríguez (2023); Martinez et al. (2024). From a physical perspective, it is well established that non-Abelian interactions exhibit a richer and more intricate phenomenology compared to their Abelian counterparts. One striking difference lies in the behaviour of their mediators—while non-Abelian gauge bosons, such as those in the nuclear forces, can interact with one another, Abelian interactions like electromagnetism lack this self-interaction property Yang and Mills (1954); Peskin and Schroeder (1995); Schwartz (2014); Weinberg (2013); ’t Hooft (2005). This fundamental distinction makes it both challenging and intriguing to uncover novel phenomena that arise uniquely from the non-Abelian nature of vector fields.
@@ II The theory
Introducing the gauge covariant derivative $D_{\alpha}F_{a\mu}^{\alpha}\equiv\nabla_{\alpha}F_{a\mu}^{\alpha}+\tilde{g}B^{b\alpha}\epsilon_{abc}F^{c}_{\mu\alpha}$, the equation of motion can be written as
$D_{\alpha}F_{a\mu}^{\alpha}=-\mu^{2}B_{a\mu}+J_{a\mu}\,,$ (2)
with $J_{a\mu}=\chi_{1}B_{a\mu}B_{b\nu}B^{b\nu}+\chi_{2}B_{a}^{\alpha}B^{b}_{\mu}B_{b\alpha}$. Due to the symmetry of the field strength tensor, the condition $D^{\mu}D_{\alpha}F_{a\mu}^{\alpha}=0$ holds. Consequently, one can properly define conserved charges [^1] There are also topological charges in this theory Martinez et al. (2023); Gómez and Rodríguez (2023). in this theory since the action remains invariant under the SU(2) group of global transformations Fuchs and Schweigert (2003); Ramond (2010). The structure of the physical charges ultimately depends on the vector field configuration. For the chosen profile in this work, these charges trivially vanish, as demonstrated in Appendix A.
@@
where $z_{ab}\equiv\bar{\mu}^{2}g_{ab}+2\chi_{2}X_{ab}$, $\bar{\mu}^{2}\equiv\mu^{2}+2\chi_{1}X$ is the dynamical mass, $X\equiv-\frac{1}{2}B_{c\nu}B^{c\nu}$, $X_{ab}\equiv-\frac{1}{2}B_{a}^{\mu}B_{b\mu}$ and $g_{ab}=\delta_{ab}$. The Lorenz condition is generally not satisfied for an arbitrary profile. However, as shown in Appendix A.1, the specific configuration employed in this work always satisfies this condition.
@@ II.1 Spherical symmetry case: the ’t Hooft-Polyakov configuration
Another notable configuration is the Dyon (see Appendix A.1), characterized by $A_{1}=\phi_{1}=0$ and $A_{0}\neq 0$, $\phi_{2}\neq 0$ (see, e.g., (Julia and Zee, 1975; Bizon and Popp, 1992; Bjoraker and Hosotani, 2000)). This configuration will not be discussed here and is left for future work. Furthermore, we will refer to the vector gauge fields $B_{\mu}^{a}$ simply as the vector field $w(t,r)$, since it represents the sole dynamical degree of freedom.
Finally, in geometric units, $w$ is dimensionless and the gauge coupling constant $\tilde{g}$ has dimensions $[\tilde{g}]=[L]^{-1}$. Consequently, we will use $\tilde{g}$ to define dimensionless variables. From this point onward, all quantities will be dimensionless, given by the following transformations:
@@ III Evolution equations in Polar-areal coordinates
$r^{2}\ddot{w}+r^{2}\dot{w}\left(\dot{B}-\dot{A}\right)-r^{2}e^{2(A-B)}\left[w^{\prime}\left(A^{\prime}-B^{\prime}\right)+w^{\prime\prime}\right]\\ +e^{2A}\left[w\left(w^{2}-1\right)+\mu^{2}r^{2}(w+1)-\chi(w+1)^{3}\right]=0.$ (12)
where overdot and a prime denote time and spatial derivatives, respectively. For simplification, we have introduced a new parameter defined as a linear combination of the self-interaction parameters: $\chi=2\chi_{1}+\chi_{2}$. This reduction leaves us with only two free parameters: the particle mass $\mu$ and the effective self-interaction parameter $\chi$.
@@
$C_{B}\equiv\partial_{r}B-\frac{Q^{2}}{r}-\frac{P^{2}}{r}-\frac{(1-e^{2B})}{2r}-\frac{e^{2B}}{2r^{3}}[(1-w^{2})^{2}\\ +2\mu^{2}r^{2}(w+1)^{2}-\chi(w+1)^{4}]=0.$ (18)
@@
Exact values of quantities $\mathcal{A}$$\mathcal{B}$,$\mathcal{Q}$, $\mathcal{R}$ and $\mathcal{S}$ can be found, for instance, in Appendix A of Franchini et al. (2022); Thaalba et al. (2024). The characteristic speeds are defined as $c\equiv-\frac{\xi_{t}}{\xi_{r}}$ and are given by
@@
Hence, the condition for strong hyperbolicity is fully guarantee Hilditch (2013). Having established that our system is strongly hyperbolic, a feature that implies it is also well-posed, we can now proceed to perform numerical simulations with suitable initial data and appropriate boundary conditions.
@@ III.1 Initial data
Here $a_{0}$ is the amplitude, $r_{0}$ is the grid location of the pulse and $w_{0}$ denotes the root-mean-square width of both pulses. In general, we choose the parameters such that $w(0,0)=-1$ and $w^{\prime}(0,0)=0$, ensuring consistency with Eq. (8). For the Type II family, enforcing these conditions leads to the derived relations: $a_{0}=-1-\coth{(\frac{r_{0}}{w_{0}})}$ and $b=-1+\coth{(\frac{r_{0}}{w_{0}})}$. No such conditions are imposed on the Type I family. We have used a standard nomenclature to ease readability and comparison  [^4] In this sense, we have ambiguously used the same letter, $w$, to describe both the width of the pulse and the (initial) value of the vector field. However, the former is denoted as $w_{0}$, while the latter is written as $w(0,r)$..
@@ III.2 Numerical scheme
$P(t,0)=0,\;\;Q(t,0)=0,$ (33)
as required by the constraints equations. Additionally, we impose the condition $w(t,0)=-1$. Furthermore, we impose outgoing Sommerfeld boundary conditions on the $Q$ and $P$ fields at $r_{\rm max}.$
Through the numerical evolution, we monitor the constraint equations by computing the cumulative error (residual) at every time step for the selected resolutions. As demonstrated in Appendix B, our simulations effectively converge with second-order accuracy as the resolution is increased.
@@ III.4 A look towards spherical collapse
The formation of a black hole within this framework raises several intriguing questions that warrant further investigation. Notably, this aligns with the existence of an exact static charged black hole solution predicted by this theory in a former study Gómez and Rodríguez (2023), albeit in the massless case $\mu=0$. This solution describes a Reissner-Nordström-like black hole with a non-Abelian magnetic charge determined by the self-interaction parameter. However, the stability of this exotic solution remains unexplored. It should be thoroughly examined to determine whether spherical collapse can indeed lead to the formation of such a black hole or potentially reveal other undiscovered compact objects.
@@ Appendix A Noether currents
where we have used the equation of motion, and $J^{a\mu}$ represents three currents given by
@@
A.1 Dyon and magnectic monopole configurations
@@
Appendix B Convergence test

### v2

@@ I Introduction
These issues were also identified in the generalized Proca theory Tasinato (2014); Heisenberg (2014); Allys et al. (2016); Beltran Jimenez and Heisenberg (2016); Gallego Cadavid and Rodriguez (2019), a vector-tensor formulation inspired by Horndeski theory Horndeski (1974); Deffayet et al. (2011). Thus, it seems that this issue is not exclusive to Proca fields alone, it also arises in generalizations of the Proca theory Garcia-Saenz et al. (2021); Ünlütürk et al. (2023). These findings pose a challenge to the effective field theory description of massive vector fields, particularly those based on effective description of interactions, including dark matter scenarios Holdom (1986); Graham et al. (2016); Agrawal et al. (2020); Goodsell et al. (2009); Bastero-Gil et al. (2019); Ema et al. (2019); Kolb and Long (2021); Ahmed et al. (2020); Caputo et al. (2021); Antypas et al. (2022). A potential resolution to this issue is that the effective field theory description breaks down before the ghost instability develops Aoki and Minamitsuji (2022). Another approach consists in fixing the equations Barausse et al. (2022); Rubio et al. (2024), a procedure that “corrects” the hyperbolicity character of the system Cayuso et al. (2017). See also Ref. Coates and Ramazanoğlu (2023c) for a discussion on the shortcomings of this approach.
@@
At this point, the reader may wonder why we focus specifically on non-Abelian SU(2) fields. The motivation stems from their potential to drive various phenomena in both cosmology Maleknejad and Sheikh-Jabbari (2013, 2011); Adshead and Wyman (2012); Garnica et al. (2022); Guarnizo et al. (2020); Garcia-Serna et al. (2025) and astrophysics Bartnik and Mckinnon (1988); Bizon (1990); Greene et al. (1993); Dzhunushaliev and Folomeev (2020); da Rocha (2020); Jain (2022); Martinez et al. (2023); Gómez and Rodr´ıguez (2023); Martinez et al. (2024). From a physical perspective, it is well established that non-Abelian interactions exhibit a richer and more intricate phenomenology compared to their Abelian counterparts. One striking difference lies in the behaviour of their mediators—while non-Abelian gauge bosons, such as those in the nuclear forces, can interact with one another, Abelian interactions like electromagnetism lack this self-interaction property Yang and Mills (1954); Peskin and Schroeder (1995); Schwartz (2014); Weinberg (2013); ’t Hooft (2005). This fundamental distinction makes it both challenging and intriguing to uncover novel phenomena that arise uniquely from the non-Abelian nature of vector fields.
@@ II The theory
Introducing the gauge covariant derivative $D_{\alpha}F_{a\mu}^{\alpha}\equiv\nabla_{\alpha}F_{a\mu}^{\alpha}+\tilde{g}B^{b\alpha}\epsilon_{abc}F^{c}_{\alpha\mu}$, the equation of motion can be written as
$D_{\alpha}F_{a\mu}^{\alpha}=\mu^{2}B_{a\mu}-J_{a\mu}\,$, (2)
with $J_{a\mu}=\chi_{1}B_{a\mu}B_{b\nu}B^{b\nu}+\chi_{2}B_{a}^{\alpha}B^{b}_{\mu}B_{b\alpha}$. Due to the symmetry of the field strength tensor, the condition $D^{\mu}D_{\alpha}F_{a\mu}^{\alpha}=0$ holds. Consequently, one can properly define conserved charges [^1] There are also topological charges in this theory Martinez et al. (2023); Gómez and Rodr´ıguez (2023). in this theory since the action remains invariant under the SU(2) group of global transformations Fuchs and Schweigert (2003); Ramond (2010). The structure of the physical charges ultimately depends on the vector field configuration. For the chosen profile in this work, these charges trivially vanish, as demonstrated in Appendix A.
@@
where $z_{ab}\equiv\bar{\mu}^{2}g_{ab}+2\chi_{2}X_{ab}$, $\bar{\mu}^{2}\equiv\mu^{2}+2\chi_{1}X$ is the dynamical mass, $X\equiv-\frac{1}{2}B_{c\nu}B^{c\nu}$, $X_{ab}\equiv-\frac{1}{2}B_{a}^{\mu}B_{b\mu}$ and $g_{ab}=\delta_{ab}$. The Lorenz condition is generally not satisfied for an arbitrary profile. However, as shown in Appendix B, the specific configuration employed in this work always satisfies this condition.
@@ II.1 Spherical symmetry case: the ’t Hooft-Polyakov configuration
Another notable configuration is the Dyon (see Appendix B), characterized by $A_{1}=\phi_{1}=0$ and $A_{0}\neq 0$, $\phi_{2}\neq 0$ (see, e.g., (Julia and Zee, 1975)). In the pure Einstein-Yang-Mills theory, the only known static, spherically symmetric, and asymptotically flat dyonic solution is the Reissner-Nordström solution with both electric and magnetic charge Galtsov and Ershov (1989); Ershov and Galtsov (1990); Bizon and Popp (1992). However, to date, no other dyon solution has been found in theories involving non-Abelian SU(2) gauge fields that are neither coupled to additional matter fields nor non-minimally coupled to gravity. This case will not be discussed further and is left for future work. Furthermore, we will refer to the vector gauge fields $B_{\mu}^{a}$ simply as the vector field $w(t,r)$, since it represents the sole dynamical degree of freedom.
Finally, in geometric units, $w$ is dimensionless and the gauge coupling constant $\tilde{g}$ has dimensions $\tilde{g}]=[L]^{-1}$. Consequently, we will use $\tilde{g}$ to define dimensionless variables. From this point onward, all quantities will be dimensionless, given by the following transformations:
@@ III Evolution equations in Polar-areal coordinates
$r^{2}\ddot{w}+r^{2}\dot{w}\left(\dot{B}-\dot{A}\right)-r^{2}e^{2(A-B)}\left[w^{\prime}\left(A^{\prime}-B^{\prime}\right)+w^{\prime\prime}\right]\\ +e^{2A}\left[w\left(w^{2}-1\right)+\mu^{2}r^{2}(w+1)-\chi(w+1)^{3}\right]=0,$ (12)
where overdot and a prime denote time and spatial derivatives, respectively. It is convenient to introduce an effective parameter, since the self-interaction parameters of the theory can be recast as a linear combination in the field equations [^3] This simplification is only possible under the ’t Hooft–Polyakov monopole configuration, even at the perturbative level. When additional degrees of freedom are present, it no longer holds, and each self-interaction term must be analyzed independently.: $\chi=2\chi_{1}+\chi_{2}$. This reduction leaves us with only two free parameters: the particle mass $\mu$ and the effective self-interaction parameter $\chi$. This implies that, in the limit $\chi\to 0$, or equivalently $\chi_{2}=-2\chi_{1}$, the present theory reduces to the standard (non-Abelian) Proca theory.
@@
$C_{B}\equiv\partial_{r}B-\frac{Q^{2}}{r}-\frac{P^{2}}{r}-\frac{(1-e^{2B})}{2r}-\frac{e^{2B}}{2r^{3}}[(1-w^{2})^{2}\\ +2\mu^{2}r^{2}(w+1)^{2}-\chi(w+1)^{4}]=0,$ (18)
@@
Exact values of quantities $\mathcal{A}$$\mathcal{B}$$\mathcal{Q}$, $\mathcal{R}$ and $\mathcal{S}$ can be found, for instance, in Appendix A of Franchini et al. (2022); Thaalba et al. (2024). The characteristic speeds are defined as $c\equiv-\frac{\xi_{t}}{\xi_{r}}$ and are given by
@@
Hence, the condition for strong hyperbolicity is fully guaranteed Hilditch (2013). Having established that our system is strongly hyperbolic, a feature that implies it is also well-posed, we can now proceed to perform numerical simulations with suitable initial data and appropriate boundary conditions.
@@ III.1 Initial data
Here $a_{0}$ is the amplitude, $r_{0}$ is the grid location of the pulse and $w_{0}$ denotes the root-mean-square width of both pulses. In general, we choose the parameters such that w(0,0)=1w(0,0)=-1 and $w^{\prime}(0,0)=0$, ensuring consistency with Eq. (8). For the Type II family, enforcing these conditions leads to the derived relations: $a_{0}=-1-\coth{(\frac{r_{0}}{w_{0}})}$ and $b=-1+\coth{(\frac{r_{0}}{w_{0}})}$. No such conditions are imposed on the Type I family. We have used a standard nomenclature to ease readability and comparison  [^5] In this sense, we have ambiguously used the same letter, $w$, to describe both the width of the pulse and the (initial) value of the vector field. However, the former is denoted as $w_{0}$, while the latter is written as $w(0,r)$..
@@ III.2 Numerical scheme
P(t,0)=0,Q(t,0)=0,P(t,0)=0,\;\;Q(t,0)=0, (33)
as required by the constraints equations. Additionally, we impose the condition w(t,0)=1w(t,0)=-1. Furthermore, we impose outgoing Sommerfeld boundary conditions on the $Q$ and $P$ fields at $r_{\rm max}.$
Through the numerical evolution, we monitor the constraint equations by computing the cumulative error (residual) at every time step for the selected resolutions. As demonstrated in Appendix D, our simulations effectively converge with second-order accuracy as the resolution is increased.
@@ III.4 A look towards spherical collapse
The formation of a black hole within this framework raises several intriguing questions that warrant further investigation. Notably, this aligns with the existence of an exact static charged black hole solution predicted by this theory in a former study Gómez and Rodr´ıguez (2023), albeit in the massless case $\mu=0$. This solution describes a Reissner-Nordström-like black hole with a non-Abelian magnetic charge determined by the self-interaction parameter. However, the stability of this exotic solution remains unexplored. It should be thoroughly examined to determine whether spherical collapse can indeed lead to the formation of such a black hole or potentially reveal other undiscovered compact objects.
III.5 Comparison with the Abelian case
While we do not provide detailed calculations for the Abelian case, we highlight the key differences between the Abelian and non-Abelian cases, which lead to distinct features concerning the well-posedness of the Cauchy problem. One crucial point of distinction lies in the configuration of the vector field. In the Abelian case, it is well known that, under spherical symmetry, only the temporal and radial components of the vector field survive and must depend solely on the temporal and radial coordinates. As a result, the vector field is manifestly invariant under rotations—an invariance that excludes the possibility of the magnetic part.
In contrast, in the SU(2) non-Abelian case, a more general configuration for the vector field is allowed within spherical symmetry, as described by Eq. (6). A particularly relevant ansatz involves a purely magnetic gauge field configuration, which, as previously noted, has no analogue in the Abelian case. Hence, direct comparisons between the two cases can be misleading, as the Abelian configuration is purely longitudinal, whereas the non-Abelian configuration is purely transverse. Nevertheless, the structure of the wave equation in both cases can be examined to identify regimes where the system transitions from hyperbolic to parabolic or elliptic behavior. Such behavior arises only in the Abelian case. See Appendix C.
Another significant aspect concerns the Lorenz constraint. In the purely magnetic non-Abelian configuration, this condition is automatically satisfied without the need for additional constraints. By contrast, in the Abelian case, the Lorenz condition must be explicitly enforced at each point in the computational grid, modifying the principal part of the wave equation (see, e.g., Coates and Ramazanoğlu (2022)). This enforcement compromises the hyperbolic character of the field equations during nonlinear evolution. Technical analyses further show that the breakdown in well-posedness arises when the system transitions to a parabolic regime or when characteristic speeds become unbounded Rubio et al. (2024). These pathologies are associated with mixed-type partial differential equations, such as Tricomi and Keldysh types Barausse et al. (2022); Ripley (2022).
Preliminary calculations for the dyon configuration (see Appendix B), which introduces an additional degree of freedom in the theory, suggest that similar dynamical shortcomings may emerge during nonlinear evolution, akin to those observed in the Abelian case. This highlights an important point: generality does not necessarily imply physical viability. Such extensions, while theoretically permissible, require dedicated follow-up studies within the present framework.
@@ Appendix A Noether currents
where we have used the field equations, and $J^{a\mu}$ represents three currents given by
@@
Appendix B Dyon and magnectic monopole configurations
@@
Appendix C Comparison between the Abelian and non-Abelian cases
As already mentioned, the key difference stems from the generalized Lorentz gauge condition, which is not automatically satisfied in the Abelian case. As a result, this imposes constraints on the initial conditions. Moreover, the generalized Lorentz gauge modifies the principal symbol. In contrast, for the non-Abelian pure magnetic monopole ansatz, no such constraint on the initial conditions arises, meaning that the principal symbol remains unmodified. To illustrate this point, we perform the analysis on flat spacetime for both models with a quartic self-interaction.
The Abelian field will be denoted by $A_{\mu}$, and the field equations are given by
$\nabla_{\mu}F^{\mu\nu}=z\mu^{2}A^{\nu},$ (42)
where $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$ and $z=(1+\lambda A_{\alpha}A^{\alpha})$.
The generalized Lorentz gauge is
$\nabla_{\mu}(zA^{\mu})=0.$ (43)
This expression is a consequence of the field equations, so it does not represent a gauge choice despite its name. It can be used to derive an alternative form of the field equations Coates and Ramazanoğlu (2022, 2023a):
$g_{\alpha\beta}^{\rm eff}\nabla^{\alpha}\nabla^{\beta}A_{\nu}+2\lambda A^{\mu}A^{\alpha}\nabla_{\mu}F_{\nu\alpha}\\ +2z\lambda\nabla_{\mu}A_{\alpha}\nabla_{\nu}(z^{-1}A^{\mu}A^{\alpha})-z^{2}\mu^{2}A_{\nu}=0,$ (44)
where $g_{\alpha\beta}^{\rm eff}=zg_{\alpha\beta}+2\lambda A_{\alpha}A_{\beta}$ is the effective metric. The field equations (42) are equivalent to (44) together with the generalized Lorentz gauge (43). For the 1+1 case, the highest derivatives come only from the term proportional to the effective metric. The presence of the disformal term in the effective metric can cause a loss of hyperbolicity, which is the central issue we aim to highlight here.
In the 1+1 spherically symmetric configuration that involves only the radial component $A_{\mu}=(0,r^{-1}v,0,0)$, the determinant of the effective metric is
$\det\Bigl{(}g_{\alpha\beta}^{\rm eff}\Bigr{)}=-\frac{\left(r^{2}+\lambda v^{2}\right)\left(r^{2}+3\lambda v^{2}\right)}{r^{4}},$ (45)
and it can vanish when $\lambda<0$, more exactly, when the field reaches the values $v_{1,2}=\pm r/(-\lambda)^{1/2},\pm r/(-3\lambda)^{1/2}$. Therefore, self-interaction can induce hyperbolicity loss. To make this more explicit, the dynamical field equation in this 1+1 case reduces to
$(1+\lambda v^{2}/r^{2})\partial_{t}^{2}v+(1+3\lambda v^{2}/r^{2})\partial_{r}^{2}v\\ +\frac{4\lambda v(\partial_{r}v)^{2}}{\lambda v^{2}+r^{2}}-\frac{4\lambda v^{2}(\lambda v^{2}+3r^{2})\partial_{r}v}{r^{3}(\lambda v^{2}+r^{2})}\\ +\frac{2(\lambda^{2}v^{5}+2\lambda r^{2}v^{3}-r^{4}v)}{r^{4}(\lambda v^{2}+r^{2})}-\frac{\mu^{2}v(\lambda v^{2}+r^{2})^{2}}{r^{4}}=0,$ (46)
where two distinctive features arising from self-interaction become evident: a modification of the principal part of the equations and nonlinearities. The former is especially critical: a wrong sign in the principal part can spoil the hyperbolic character of the system, causing frequency modes to grow exponentially over time. As a result, even small initial perturbations may evolve into large amplitudes, resulting in a breakdown of predictivity.
In contrast, for the non-Abelian case under the ’t Hooft–Polyakov configuration, the field equation takes the form
$\partial_{t}^{2}w+\partial_{r}^{2}w-\frac{w(w^{2}-1)}{r^{2}}-\mu^{2}(w+1)+\frac{\chi(w+1)^{3}}{r^{2}}=0.$ (47)
This equation remains hyperbolic regardless of the value of the self-interaction parameter $\chi$, and no additional constraints on the initial data are required.
Appendix D Convergence test